### v1

# Lecture Notes on High Dimensional Linear Regression

Lecture Notes on High Dimensional Linear Regression
Alberto Quaini111 Department of Econometrics, Erasmus University of Rotterdam, P.O. Box 1738, 3000 DR Rotterdam, The Netherlands. 222 Please, let me know if you find typos or mistakes at quaini@ese.eur.nl.
(December 20, 2024 (first version: October 17, 2024))
Introduction
These lecture notes were developed for a Master’s course in advanced machine learning at Erasmus University of Rotterdam. The course is designed for graduate students in mathematics, statistics and econometrics. The content follows a proposition-proof structure, making it suitable for students seeking a formal and rigorous understanding of the statistical theory underlying machine learning methods.
At present, the notes focus on linear regression, with an in-depth exploration of the existence, uniqueness, relations, computation, and non-asymptotic properties of the most prominent estimators in this setting: least squares, ridgeless, ridge, and lasso.
Background
It is assumed that readers have a solid background in calculus, linear algebra, convex analysis, and probability theory. Some definitions and results from these fields, relevant for the course, are provided in the Appendix for reference.
Book-length references
The content of these lecture notes is inspired by a wide range of existing literature, but the presentation of topics follows my own interpretation and logical structure. Although most of the content can be traced back to established sources, certain sections reflect my perspective, and some material is original to this course. For those interested in more comprehensive, book-length discussions of related topics, the following key references are recommended: Hastie et al. (2009), Bühlmann and Van De Geer (2011), Hastie et al. (2015), and Wainwright (2019).
Disclaimer
Please note that despite my efforts, these lecture notes may contain errors. I welcome any feedback, corrections, or suggestions you may have. If you spot any mistakes or have ideas for improvement, feel free to contact me via email at quaini@ese.eur.nl.
Notation
  • All random variables are defined on complete probability space $(\Omega,\mathcal{F},\mathbb{P})$ and take values in a real Euclidean space.
  • For a random variable (vector) [matrix] $x$ ($\bm{x}$) [$\bm{X}$], the notation $x\in\mathbb{R}$ ($\bm{x}\in\mathbb{R}^{n}$) [$\bm{X}\in\mathbb{R}^{n\times p}$] means that $x$ ($\bm{x}$) [$\bm{X}$] takes values in $\mathbb{R}$ ($\mathbb{R}^{n}$) [$\mathbb{R}^{n\times p}$].
[…]
  • Given a random variable $x$, its expectation is denoted $\mathbb{E}[x]$ and its variance $\operatorname{\mathbb{V}ar}[x]$.
  • For a vector $\bm{x}\in\mathbb{R}^{n}$, the $i-$th element is denoted $x_{i}$ for $i=1,\ldots,n$.
  • For a matrix $\bm{A}\in\mathbb{R}^{n\times p}$, the $i,j-$th element is denoted $A_{i,j}$, the $j-$th column is denoted $\bm{A}_{j}$ and the $i-$th row is denoted $A_{(i)}$, for $i=1,\ldots,n$ and $j=1,\ldots,p$.
  • The transpose of a matrix $\bm{A}\in\mathbb{R}^{n\times p}$ is denoted $\bm{A}^{\prime}$.
  • The Moore-Penrose inverse of a matrix $\bm{A}\in\mathbb{R}^{n\times p}$ is denoted $\bm{A}^{+}$.
  • The $l_{p}-$norm $\left\lVert\cdot\right\rVert_{p}$ on $\mathbb{R}^{n}$ is defined for all $\bm{v}\in\mathbb{R}^{n}$ as $\left\lVert\bm{v}\right\rVert_{p}:=\left(\sum_{i=1}^{n}|v_{i}|^{p}\right)^{1/p}$ when $p\in[1,+\infty)$, and $\left\lVert\bm{v}\right\rVert_{p}:=\max_{i=1}^{n}|v_{i}|$ when $p=+\infty$.
  • Given vector $\bm{\theta}\in\mathbb{R}^{p}$, the $l_{0}-$norm (which is not a norm!) $\left\lVert\bm{x}\right\rVert_{0}$ counts the number of nonzero elements of $\bm{\theta}$.
  • $\operatorname*{argmin}_{x\in X}f(x)$ denotes the set of minimizers of $f$ over set $X$.
  • $\operatorname{diag}(\bm{A})$ denotes the diagonal elements of a square matrix $\bm{A}\in\mathbb{R}^{n\times n}$
  • $\operatorname{diag}(a_{1},\ldots,a_{n})$ denotes a square matrix in $\mathbb{R}^{n\times n}$ that has diagonal elements given by $a_{1},\ldots,a_{n}\in\mathbb{R}$ and that has zero elsewhere.
  • Given a matrix $\bm{A}\in\mathbb{R}^{n\times p}$, its rank is $\operatorname{Rank}(\bm{A})$, its range is $\operatorname{Range}(\bm{A})$, its kernel is $\operatorname{Ker}(\bm{A})$, its trace is denoted $\operatorname{Trace}(\bm{A})$.
  • We denote $\operatorname{P}_{S}$ the orthogonal projection onto set $S\subset\mathbb{R}^{n\times p}$.
  • Given a vector space $V$, the sum of two subsets $A,B\subset V$ is defined as $A+B:=\{a+b\ :\ a\in A,b\in B\}$. The sum of a set $A\in V$ and a vector $b\in V$ is defined as $A+b:=\{a+b\ :\ a\in A\}$.
  • The symbol $\partial$ indicates the subdifferential.
Chapter 1 Linear Regression
Linear regression is a supervised learning technique aimed at predicting a target random variable $y$ using a linear combination
$\bm{x}^{\prime}\bm{\theta}=x_{1}\theta_{1}+\ldots+x_{p}\theta_{p}$
of explanatory variables $\bm{x}=[x_{1},\ldots,x_{p}]^{\prime}$, where $\bm{\theta}\in\mathbb{R}^{p}$ and $p\in\mathbb{N}$111 An intercept in $f(\bm{x};\bm{\theta})$ can be introduced by adding a constant term to the predictors. The target variable $y$ is also referred to as dependent or output variable, while the explanatory variables $\bm{x}$ are also known as independent variables, predictors or input variables.
In typical applications, we observe only a sample of size $n\in\mathbb{N}$ of these random variables, represented by the pairs $(\bm{x}_{i},y_{i})_{i=1}^{n}$, where $\bm{x}_{i}\in\mathbb{R}^{p}$ and $y_{i}\in\mathbb{R}$ for each $i$. Given a regression coefficient $\bm{\theta}_{0}\in\mathbb{R}^{p}$, a statistical linear model, or simply linear model, is expressed as
$y_{i}=\bm{x}_{i}^{\prime}\bm{\theta}_{0}+\varepsilon_{0i},\quad i=1,\ldots,n,$ (1.1)
where $\varepsilon_{0i}$ are real-valued residual random variables. Figure 1.1 depicts a linear model for $i=1,\ldots,n$ observations $y_{i}$, with two predictors $\tilde{\bm{x}}_{i}=[1,x_{i}]^{\prime}$ consisting of a unit constant and a variable $x_{i}$, a coefficient $\bm{\theta}_{0}\in\mathbb{R}^{2}$ and the associated error terms $\varepsilon_{0i}$. The Data Generating Process (DGP), i.e., the joint distribution of the predictors $\bm{x}$ and the real-valued residual random variables $\bm{\varepsilon}_{0}=[\varepsilon_{0i}]_{i=1}^{n}$, is subject to certain restrictions. Depending on the type of restrictions imposed on the DGP, different types of linear models are obtained. The two general forms of linear models are fixed and random design models, which are defined as follows.
Definition 1 (Fixed design model).
In a fixed design model, the sequence $(\bm{x}_{i})_{i=1}^{n}$ is fixed. The residuals $\varepsilon_{0i}$ are independent and identically distributed.
Definition 2 (Random design model).
In a random design model, the pair $(\bm{x}_{i},y_{i})_{i=1}^{n}$ is a sequence of independent and identically distributed random variables.
The fixed design model is particularly suitable when the predictors are controlled by the analyst, such as the dose of medication administered to patients in the treatment group in a clinical trial. Conversely, the random design model is appropriate when the explanatory variables are stochastic, such as the wind speed observed at a specific time and location.
Figure 1.1: Statistical linear model $y_{i}=\tilde{\bm{x}}_{i}^{\prime}\bm{\theta}_{0}+\varepsilon_{0i}$ where $\tilde{\bm{x}}_{i}=[1,x_{i}]^{\prime}$ and $\bm{\theta}_{0}=[0.5,0.8]^{\prime}$.
We organize the observed values of the target variable in the vector $\bm{y}=[y_{1},\ldots,y_{n}]^{\prime}\in\mathbb{R}^{n}$, and the observations on the predictors in the design matrix
$\bm{X}:=\begin{bmatrix}x_{11}&\ldots&x_{1p}\\ \vdots&\ddots&\vdots\\ x_{n1}&\ldots&x_{np}\end{bmatrix}\mathbb{R}^{n\times p}.$
With this notation, the linear model (1.1) can be expressed as:
$\bm{y}=\bm{X}\bm{\theta}_{0}+\bm{\varepsilon}_{0},$
where $\bm{\varepsilon}_{0}=[\varepsilon_{01},\ldots,\varepsilon_{0n}]^{\prime}\in\mathbb{R}^{n}$.
Example 1.
A classic example of linear regression is found in the work of Nerlove et al. (1961), which examines returns to scale in the U.S. electricity power supply industry. In this study, the total cost $y_{i}$ for firm $i$ is predicted using a linear model based on the firm’s output production $x_{i1}$, the wage rate $x_{i2}$, the price of fuel $x_{i3}$, and the rental price of capital $x_{i4}$, with data from $n=145$ electric utility companies.
The rest of the chapter is organized as follows. First, we study the most basic linear regression approach, the method of least squares projection, and some of its recent machine learning extensions. Our study focuses on their existence, uniqueness, connections, geometric interpretation, and computation. Then, we study we cover both their finite- or small-sample properties, that are valid for any given sample size, and their asymptotic properties, that are useful approximations when the sample size is large enough.
1.0.1 Estimators and their properties
Definition 3 (Estimand).
An estimand is a feature, or parameter, of interest of the population.
Definition 4 (estimator and estimate).
An estimator is a function taking as input the data, and possibly other auxiliary variables, and outputting an estimate, which is a specific value assigned to the estimand.
For instance, in the context of the linear model (1.1), the coefficient $\bm{\theta}_{0}\in\mathbb{R}^{p}$ represents the estimand. An estimator is a function $\hat{\bm{\theta}}_{n}:\mathbb{R}^{n}\times\mathbb{R}^{n\times p}\to\mathbb{R}^{p}$ that takes as inputs the data $(\bm{y},\bm{X})\in\mathbb{R}^{n}\times\mathbb{R}^{n\times p}$ and produces an estimate $\hat{\bm{\theta}}_{n}(\bm{y},\bm{X})\in\mathbb{R}^{p}$. For simplicity, we use the same notation, $\hat{\bm{\theta}}_{n}$, to refer to both the estimator and the resulting estimate, although formally the estimate should be written as $\hat{\bm{\theta}}_{n}(\bm{y},\bm{X})$. We may also write $\hat{\bm{\theta}}_{n}\in\mathbb{R}^{p}$ to indicate that an estimator outputs values in $\mathbb{R}^{p}$.
Definition 5 (Linear prediction).
The quantity $\operatorname{lm}(\bm{X},\bm{\theta}):=\bm{X}\bm{\theta}$ denotes the linear prediction associated to coefficient vector $\bm{\theta}\in\mathbb{R}^{p}$ and predictors $\bm{X}\in\mathbb{R}^{n\times p}$.
Let $\bm{M}^{+}$ denote the Moore-Penrose inverse of a generic real-valued matrix $\bm{M}$. We make extensive use of the following useful projections.222 We use the notation $\operatorname{Range}$ ($\operatorname{Ker}$) for the range (kernel) of a matrix. Details on these sets, the Moore-Penrose inverse and orthogonal projections are given in Appendix Section LABEL:app:sec:linear_algebra.
Definition 6 (Useful projections).
Given a fixed matrix $\bm{X}\in\mathbb{R}^{n\times p}$:
  • $\bm{\operatorname{P}}_{\operatorname{Range}(\bm{X}^{\prime})}:=\bm{X}^{+}\bm{X}$ is the orthogonal projector onto $\operatorname{Range}(\bm{X}^{\prime})$;
  • $\bm{\operatorname{P}}_{\operatorname{Ker}(\bm{X})}:=\bm{I}-\bm{X}^{+}\bm{X}$ is the orthogonal projector onto $\operatorname{Ker}(\bm{X})$;
  • $\bm{\operatorname{P}}_{\operatorname{Range}(\bm{X})}:=\bm{X}\bm{X}^{+}$ is the orthogonal projector onto $\operatorname{Range}(\bm{X})$;
  • $\bm{\operatorname{P}}_{\operatorname{Ker}(\bm{X}^{\prime})}:=\bm{I}-\bm{X}\bm{X}^{+}$ is the orthogonal projector onto $\operatorname{Ker}(\bm{X}^{\prime})$.
The next proposition demonstrates that, if we fix the design matrix $\bm{X}$, we can focus on regression coefficients in $\operatorname{Range}(\bm{X}^{\prime})$. Indeed, coefficients in this set span all possible linear predictions that can be achieved through $\bm{X}$.
Proposition 1.0.1.
Given a matrix $\bm{X}\in\mathbb{R}^{n\times p}$, for any $\bm{\theta}\in\mathbb{R}^{p}$:
$\operatorname{lm}(\bm{X},\bm{\theta})=\operatorname{lm}(\bm{X},\bm{\operatorname{P}}_{\operatorname{Range}(\bm{X}^{\prime})}\bm{\theta}).$
Proof.
Using the identity $\bm{I}=\bm{\operatorname{P}}_{\operatorname{Range}(\bm{X}^{\prime})}+\bm{\operatorname{P}}_{\operatorname{Ker}(\bm{X})}$, we have for any $\bm{\theta}\in\mathbb{R}^{p}$:
$\bm{X}\bm{\theta}=\bm{X}(\bm{\operatorname{P}}_{\operatorname{Range}(\bm{X}^{\prime})}+\bm{\operatorname{P}}_{\operatorname{Ker}(\bm{X})})\bm{\theta}=\bm{X}\bm{\operatorname{P}}_{\operatorname{Range}(\bm{X}^{\prime})}\bm{\theta}.$
Finite-sample properties
Since an estimator is derived from data, it is a random variable. Intuitively, when comparing two estimators of the same estimand, we prefer the one whose probability distribution is ”more concentrated” around the true value of the estimand. Formally, estimators are compared using several key properties.333 Some of these properties are defined by means of the $l_{2}-$norm. Note that this choice is typical, but arbitrary.
Definition 7 (Bias).
The bias of an estimator $\hat{\bm{\theta}}_{n}$ of $\bm{\theta}_{0}$ is the difference between the expected value of the estimator and the estimand:
$\operatorname{Bias}(\hat{\bm{\theta}}_{n},\bm{\theta}_{0}):=\mathbb{E}[\hat{\bm{\theta}}_{n}]-\bm{\theta}_{0}.$
Definition 8 (Estimation risk).
The estimation risk of an estimator $\hat{\bm{\theta}}_{n}$ of $\bm{\theta}_{0}$ measures the difference between the estimate and the estimand as:
$\operatorname{ER}(\hat{\bm{\theta}}_{n},\bm{\theta}_{0}):=\lVert\hat{\bm{\theta}}_{n}-\bm{\theta}_{0}\rVert_{2}^{2}.$
Definition 9 (MSE).
The Mean Squared Error (MSE) of an estimator $\hat{\bm{\theta}}_{n}$ of $\bm{\theta}_{0}$ is the expected estimation risk of the estimator:
$\operatorname{MSE}(\hat{\bm{\theta}}_{n},\bm{\theta}_{0}):=\mathbb{E}[\lVert\hat{\bm{\theta}}_{n}-\bm{\theta}_{0}\rVert_{2}^{2}].$
Definition 10 (Predictive risk).
The predictive risk of an estimator $\hat{\bm{\theta}}_{n}$ of $\bm{\theta}_{0}$ measures the difference between the linear predictions of $\hat{\bm{\theta}}_{n}$ and those of $\bm{\theta}_{0}$:
$\operatorname{PR}(\hat{\bm{\theta}}_{n},\bm{\theta}_{0}):=\lVert\operatorname{lm}(\bm{X},\hat{\bm{\theta}}_{n})-\operatorname{lm}(\bm{X},\bm{\theta}_{0})\rVert_{2}^{2}/n.$
Definition 11 (Mean predictive risk).
The Mean Predictive Risk (MPR) of an estimator $\hat{\bm{\theta}}_{n}$ of $\bm{\theta}_{0}$ is the expected predictive risk of the estimator:
$\operatorname{MPR}(\hat{\bm{\theta}}_{n},\bm{\theta}_{0}):=\mathbb{E}[\lVert\operatorname{lm}(\bm{X},\hat{\bm{\theta}}_{n})-\operatorname{lm}(\bm{X},\bm{\theta}_{0})\rVert_{2}^{2}/n].$
As a corollary to Proposition 1.0.1, the predictive risk of an estimator is unchanged if both the estimator and the estimand are projected onto $\operatorname{Range}(\bm{X}^{\prime})$.
Corollary 1.0.1.
Given a matrix $\bm{X}\in\mathbb{R}^{n\times p}$, for any estimator $\hat{\bm{\theta}}_{n}$ of $\bm{\theta}_{0}\in\mathbb{R}^{p}$:
$\operatorname{PR}(\hat{\bm{\theta}}_{n},\bm{\theta}_{0})=\operatorname{PR}(\bm{\operatorname{P}}_{\operatorname{Range}(\bm{X}^{\prime})}\hat{\bm{\theta}}_{n},\bm{\operatorname{P}}_{\operatorname{Range}(\bm{X}^{\prime})}\bm{\theta}_{0}).$
Proof.
The result follows from Proposition 1.0.1. ∎
The next proposition justifies the definition of mean predictive risk given in Definition 11.
Proposition 1.0.2.
Assume that the linear model (1.1) holds with $\mathbb{E}[\bm{x}\varepsilon_{0}]=\bm{0}$. Then, for any $\bm{\theta}\in\mathbb{R}^{p}$:
$\mathbb{E}[\lVert\bm{y}-\bm{X}\bm{\theta}\rVert_{2}^{2}/n]=\operatorname{MPR}(\bm{\theta},\bm{\theta}_{0})+\mathbb{E}[\left\lVert\bm{\varepsilon}_{0}\right\rVert_{2}^{2}/n].$
Proof.
Since $\bm{y}=\bm{X}\bm{\theta}_{0}+\bm{\varepsilon}_{0}$, we have
$\mathbb{E}[\lVert\bm{y}-\bm{X}\bm{\theta}\rVert_{2}^{2}/n]=\mathbb{E}[\lVert\bm{X}(\bm{\theta}_{0}-\bm{\theta})\rVert_{2}^{2}/n]+\mathbb{E}[\left\lVert\bm{\varepsilon}_{0}\right\rVert_{2}^{2}/n]+2(\bm{\theta}_{0}-\bm{\theta})\mathbb{E}[\bm{X}^{\prime}\bm{\varepsilon}_{0}].$
Then, the result follows since $\mathbb{E}[\bm{X}^{\prime}\bm{\varepsilon}_{0}]=\sum_{i=1}^{n}\mathbb{E}[\bm{X}_{i}\varepsilon_{0i}]=\bm{0}$, where $\bm{X}_{i}$ denotes the $i-$th row of $\bm{X}$. ∎
If our primary goal is to accurately predict the target variable, we seek a estimators $\hat{\bm{\theta}}$ with a low mean prediction risk $\mathbb{E}[\lVert\bm{y}-\bm{X}\hat{\bm{\theta}}\rVert_{2}^{2}/n]$. Since we cannot control the error term $\varepsilon_{0}$, Proposition 1.0.2 suggests that we should focus on estimators with a low mean predictive risk.
On the other hand, if our interest lies in understanding which predictors influence the target variable and how they do so, the true coefficient $\bm{\theta}_{0}$ becomes our focus. In this case, we might prefer unbiased estimators – those with zero bias – over biased ones. However, estimators with lower mean squared error (MSE) are generally favored, even if they feature some bias. The following proposition demonstrates that the MSE can be decomposed into a bias and a variance term.
Proposition 1.0.3 (Bias-variance decomposition of MSE).
Given an estimator $\hat{\bm{\theta}}_{n}\in\mathbb{R}^{p}$ for $\bm{\theta}_{0}\in\mathbb{R}^{p}$, the MSE can be decomposed as follows:
$\operatorname{MSE}(\hat{\bm{\theta}}_{n},\bm{\theta}_{0})=\lVert\operatorname{Bias}(\hat{\bm{\theta}}_{n},\bm{\theta}_{0})\rVert_{2}^{2}+\operatorname{Trace}(\operatorname{\mathbb{V}ar}[\hat{\bm{\theta}}_{n}]).$
Proof.
The result follows from
$\operatorname{MSE}(\hat{\bm{\theta}}_{n},\bm{\theta}_{0})$ $=\mathbb{E}[(\hat{\bm{\theta}}_{n}-\bm{\theta}_{0})^{\prime}(\hat{\bm{\theta}}_{n}-\bm{\theta}_{0})]$
$=\mathbb{E}[\operatorname{Trace}\{(\hat{\bm{\theta}}_{n}-\bm{\theta}_{0})(\hat{\bm{\theta}}_{n}-\bm{\theta}_{0})^{\prime}\}]$
$=\operatorname{Trace}(\mathbb{E}[(\hat{\bm{\theta}}_{n}-\bm{\theta}_{0})(\hat{\bm{\theta}}_{n}-\bm{\theta}_{0})^{\prime}])$
$=\operatorname{Trace}(\mathbb{E}[(\hat{\bm{\theta}}_{n}-\mathbb{E}[\hat{\bm{\theta}}_{n}]+\operatorname{Bias}(\hat{\bm{\theta}}_{n},\bm{\theta}_{0}))(\hat{\bm{\theta}}_{n}-\mathbb{E}[\hat{\bm{\theta}}_{n}]+\operatorname{Bias}(\hat{\bm{\theta}}_{n},\bm{\theta}_{0}))^{\prime}])$
$=\operatorname{Trace}(\operatorname{\mathbb{V}ar}[\hat{\bm{\theta}}_{n}]+\operatorname{Bias}(\hat{\bm{\theta}}_{n},\bm{\theta}_{0})\operatorname{Bias}(\hat{\bm{\theta}}_{n},\bm{\theta}_{0})^{\prime}+$
$\qquad\qquad\ \mathbb{E}[\hat{\bm{\theta}}_{n}-\mathbb{E}[\hat{\bm{\theta}}_{n}]]\operatorname{Bias}(\hat{\bm{\theta}}_{n},\bm{\theta}_{0})^{\prime}+\operatorname{Bias}(\hat{\bm{\theta}}_{n},\bm{\theta}_{0})\mathbb{E}[\hat{\bm{\theta}}_{n}-\mathbb{E}[\hat{\bm{\theta}}_{n}]]^{\prime})$
$=\operatorname{Trace}(\operatorname{\mathbb{V}ar}[\hat{\bm{\theta}}_{n}]+\operatorname{Bias}(\hat{\bm{\theta}}_{n},\bm{\theta}_{0})\operatorname{Bias}(\hat{\bm{\theta}}_{n},\bm{\theta}_{0})^{\prime})$
$=\lVert\operatorname{Bias}(\hat{\bm{\theta}}_{n},\bm{\theta}_{0})\rVert_{2}^{2}+\operatorname{Trace}(\operatorname{\mathbb{V}ar}[\hat{\bm{\theta}}_{n}]).$
Loosely speaking, the bias and the variance of an estimator are linked to the estimator’s ”complexity”. Estimators with higher complexity often fit the data better, resulting in lower bias, but they are more sensitive to data variations, leading to higher variance. Conversely, estimators with lower complexity tend to have lower variance but higher bias, a phenomenon known as the bias-variance tradeoff.
Apart from simple cases, computing the finite-sample properties of estimators, such as their MSE or predictive risk, is infeasible or overly complicated. This is because they require computations under the DGP of complex transformations of the data. When direct computation is not possible, we can rely on concentration inequalities or asymptotic approximations.
Concentration inequalities are inequalities that bound the probability that a random variable deviates from a particular value, typically its expectation. In this chapter, we focus on inequalities that control the MSE or predictive risk of an estimator, such as:
$\mathbb{P}[d(\hat{\bm{\theta}}_{n},\bm{\theta}_{0})\leq h(\bm{y},\bm{X},n,p)]\geq 1-\delta,$
or
$\mathbb{P}[d(\operatorname{lm}(\bm{X},\hat{\bm{\theta}}_{n}),\operatorname{lm}(\bm{X},\bm{\theta}_{0}))\leq h(\bm{y},\bm{X},n,p)]\geq 1-\delta,$
where $\delta\in(0,1)$ is the level of confidence, $d:\mathbb{R}^{p}\times\mathbb{R}^{p}\to[0,+\infty)$ is a distance, and $h$ is a real-valued function of the data, the sample size, and the number of predictors.
Large-sample properties
Large-sample or asymptotic theory provides an alternative approach to study and analyse estimators. Classically, this framework develops approximations of the finite-sample properties of estimators, such as their distribution, MSE or predictive risk, by letting the sample size $n\to\infty$. Consequently, these approximations work well when the sample size $n$ is much larger than the number of predictors $p$. More recently, asymptotic approximations are also developed by letting $p\to\infty$, or having both $n,p\to\infty$ at some rate. Note that, given a sample of size $n$ and number of variables $p$, there is no general indication on how to choose the appropriate asymptotic regime for $n$ and $p$, as the goodness of fit of the corresponding asymptotic approximations should be assessed on a case by case basis. In this chapter, we work with two notions from large-sample theory: consistency and asymptotic distribution.
Definition 12 (Consistency).
Estimator $\hat{\bm{\theta}}_{n}$ of $\bm{\theta}_{0}$ is consistent, written $\hat{\bm{\theta}}_{n}\overset{\mathbb{P}}{\to}\bm{\theta}_{0}$ as $n\to\infty$, if for all $\varepsilon>0$,
$\lim_{n\to\infty}\mathbb{P}[\lvert\hat{\bm{\theta}}_{n}-\bm{\theta}_{0}\rvert>\varepsilon]=0.$
Definition 13 (Asymptotic distribution).
[…]